# Connectivity of partial basis complexes of freely decomposable groups

### v1

@@ Abstract.
We show that the complex of partial bases of the free group of rank $n$, where vertices are seen up to conjugation, is Cohen–Macaulay of dimension $n-1$. This positively answers a conjecture raised by Day and Putman. We prove our results in the more general context of freely decomposable groups.
@@ 1. Introduction
Denote by $F_{n}$ the free group of rank $n$. Write $[x]$ for the conjugacy class of an element or subgroup $x$ of $F_{n}$. The complex of partial bases of $F_{n}$ is the complex $\mathcal{B}_{n}$ whose maximal simplices are sets $\{C_{1},\ldots,C_{n}\}$ of conjugacy classes such that there exists a basis $\{v_{1},\ldots,v_{n}\}$ of $F_{n}$ with $C_{i}=[v_{i}]$. It is not hard to show that $\mathcal{B}_{n}$ is a pure simplicial complex of dimension $n-1$.
This complex was studied by Day and Putman in [DP:complexpartialbases]. They showed that $\mathcal{B}_{n}$ is connected for $n\geq 2$, simply connected if $n\geq 3$, and that the orbit complex $\mathcal{B}_{n}/\mathrm{IA}_{n}$ is $(n-2)$-connected, i.e. spherical of dimension $n-1$. Here, $\mathrm{IA}_{n}$ is the Torelli subgroup of $\operatorname{Aut}(F_{n})$, defined as the kernel of the map $\operatorname{Aut}(F_{n})\to\operatorname{GL}_{n}(\mathbb{Z})$. Day–Putman conjectured [DP:complexpartialbases, Conjecture 1.1] that $\mathcal{B}_{n}$ is $(n-2)$-connected as well. The aim of this note is to prove this conjecture:
@@
The simplicial complex $\mathcal{B}_{n}$ is Cohen–Macaulay of dimension $n-1$.
We prove a more general version of Theorem 1.1, building on work by the first named author and Gupta [BG:Homotopytypecomplex] and generalisations in [Bru:buildingsfreefactor]. In the present text, we will mostly refer to [Bru:buildingsfreefactora], which contains the material of both [Bru:buildingsfreefactor] and [BG:Homotopytypecomplex].
Let $A$ be a finitely generated group that is freely decomposable, i.e. is isomorphic to a non-trivial free product. A free factor system of $A$ is a finite set $\mathcal{A}=\left\{[A_{1}],\ldots,[A_{k}]\right\}$ of conjugacy classes of non-trivial subgroups $A_{i}\leq A$ such that there is a free factorisation $A=H\ast A_{1}\ast\cdots\ast A_{k}$, where $H$ is some free subgroup of $A$. We allow $\mathcal{A}$ to be empty, but note that the empty set is only a free factor system if $A$ is a free group itself. The corank of $\mathcal{A}$ is the rank of a free complement $H$, that is, $\operatorname{crk}(\mathcal{A})=\operatorname{rk}(H)$, which is well-defined by the Grushko Theorem.
@@
For $A=F_{n}$ and $\mathcal{A}=\emptyset$, we have $c=n-2$ and the result is [BG:Homotopytypecomplex, Theorem B]. In [Bru:buildingsfreefactora, Proposition 4.30], it was shown that $\mathcal{FF}(A,\mathcal{A})$ is homotopy equivalent to the simplicial boundary of relative versions of Culler–Vogtmann Outer space [CV:Moduligraphsautomorphisms] as defined by Guirardel–Levitt [GL:outerspacefree], so Theorem 1.2 implies that these boundaries are highly connected.
The frame complex associated to $(A,\mathcal{A})$, denoted by $\mathcal{PB}(A,\mathcal{A})$, is the subposet of $\mathcal{FF}(A,\mathcal{A})\cup\{[A]\}$ given by free factor systems $\mathcal{A}^{\prime}$ such that for each $[A^{\prime}]\in\mathcal{A}^{\prime}$, we either have $[A^{\prime}]\in\mathcal{A}$ or $A^{\prime}\cong\mathbb{Z}$ is a free subgroup of $A$ of rank $1$. Note that $\{[A]\}\in\mathcal{PB}(A,\mathcal{A})$ if and only if $\mathcal{A}=\emptyset$ and $A$ is a free group of rank $1$, and in such a case $\mathcal{PB}(A,\mathcal{A})=\{[A]\}$. In general, the elements of $\mathcal{PB}(A,\mathcal{A})$ are of the form $\left\{[\langle v_{1}\rangle],\ldots,[\langle v_{i}\rangle]\right\}\sqcup\mathcal{A}$, where $\{v_{1},\ldots,v_{i}\}$ is a partial basis of the free group $H$ in a decomposition $A=H\ast A_{1}\ast\cdots\ast A_{k}$. The poset relation of $\mathcal{FF}(A,\mathcal{A})$ restricts to the subset relation on $\mathcal{PB}(A,\mathcal{A})$, so we can see $\mathcal{PB}(A,\mathcal{A})$ as a simplicial complex with vertices of the form $\left\{[\langle v\rangle]\right\}\sqcup\mathcal{A}$.
@@
We deduce Theorem 1.3 from Theorem 1.2 by showing that the inclusion map $\mathcal{PB}(A,\mathcal{A})\to\mathcal{FF}(A,\mathcal{A})$ is highly connected. For $A=F_{n}$ and $\mathcal{A}=\emptyset$, we have $\operatorname{crk}(\mathcal{A})=n$ and $\mathcal{PB}(F_{n},\emptyset)$ is an inflation of the complex $\mathcal{B}_{n}$ (see [BWW:PosetFiberTheorems, Section 6]), so Theorem 1.1 is an immediate consequence of Theorem 1.3.
Related work
The complex $\mathcal{B}_{n}$ comes with a natural action of $\operatorname{Out}(F_{n})$, the outer automorphism group of the free group. More generally, relative outer automorphism groups $\operatorname{Out}(A,\mathcal{A}^{t})$ (see [Bru:buildingsfreefactora, Section 4.1.2]) act on $\mathcal{PB}(A,\mathcal{A})$ and $\mathcal{FF}(A,\mathcal{A})$. We would like to mention that variants of the complexes $\mathcal{B}_{n}$ and $\mathcal{PB}(A,\mathcal{A})$ related to several other groups are known to be highly connected as well: An $\operatorname{Aut}(F_{n})$-version of Theorem 1.1 was established by Sadofschi Costa [SC:complexOfPartialBases], who showed that the simplicial complex whose simplices are the partial bases of $F_{n}$ (without modding out by the conjugation action on vertices) is Cohen–Macaulay of dimension $n-1$. Our paper uses different techniques and is not based on his results. Furthermore, Maazen [Maa:Homologystabilitygeneral] showed that a $\operatorname{GL}_{n}$-version of $\mathcal{B}_{n}$ is Cohen–Macaulay; Harer [Har:Stabilityhomologymapping] proved that a version related to the mapping class group of a surface is highly connected (cf. the introduction of [DP:complexpartialbases]); Putman studied a version related to the integral symplectic group $\operatorname{Sp}_{2n}(\mathbb{Z})$ [Brueck2023, Put:infinitepresentationTorelli].
@@ Acknowledgements
The first author was supported by the Deutsche Forschungsgemeinschaft (Project 427320536–SFB 1442) and Germany’s Excellence Strategy grant EXC 2044–390685587. The second author was supported by the Alexander von Humboldt Stiftung, and the FWO grant 12K1223N.
@@ 2. Posets of graphs
For the proof of Theorem 1.2 in Section 3, we compare $\mathcal{FF}(A,\mathcal{A})$ with a poset of free splittings of $A$, and study connectivity of fibres. These fibres coincide with certain finite posets of graphs, which we show to be highly connected in the present section.
@@ 2.1. Posets and wedges of spheres
We call a map $f\colon X\to Y$ between two posets a poset map if $x\leq y$ implies $f(x)\leq f(y)$. If $Y$ is a subposet of $X$, we call $f\colon X\to Y$ monotone if either for all $x\in X$ we have $f(x)\leq x$, or for all $x\in X$ we have $f(x)\geq x$. By [Qui:Homotopypropertiesposet, 1.3], such a monotone poset map $f$ defines a homotopy equivalence from $X$ to its image $f(X)$.
@@ 2.2. Graphs
In what follows, by a graph we mean a $1$-dimensional compact CW-complex with no isolated vertices. For a graph $G$, we denote by $V(G)$ its set of vertices, and by $E(G)$ its set of edges. Note that multiple edges and loops are allowed. By a subgraph we mean an edge-induced subgraph. That is, subgraphs of $G$ are in one-to-one correspondence with subsets of $E(G)$. Recall that the valence of a vertex $v\in V(G)$ is the number of connected components obtained by deleting $v$ from a small enough neighbourhood of $v$, i.e. the number of half-edges adjacent at $v$. An edge that has an endpoint with valence one is called a leaf. An edge is said to be separating if removing any point of its interior disconnects the graph. In particular, we consider leaves to be separating. The rank of a connected graph $G$, denoted by $\operatorname{rk}(G)$, is the rank of its fundamental group.
A labelled graph is a pair $(G,l)$ where $G$ is a graph, $k$ is a non-negative integer, and $l:\{1,\ldots,k\}\to V(G)$ is an injective function. We call the image of $l$ the labelled vertices of $G$ and write $k(l)$ for the number of labelled vertices. A core subgraph of a labelled graph $(G,l)$ is a non-empty subgraph such that every vertex of valence one is labelled. Observe that every labelled graph $(H,l)$ contains a unique maximal core subgraph $\mathring{H}$: this is the subgraph obtained by repeatedly removing leaves with non-labelled endpoints.
Let $e\in E(G)$ be an edge of a labelled graph $(G,l)$. We write $G\setminus e$ for the subgraph of $G$ on $E(G)\setminus\{e\}$ and $G/e$ for the graph obtained from $G$ by collapsing $e$. Both $G\setminus e$ and $G/e$ have obvious labellings induced by $l$. For example, if $e$ is a leaf with a labelled endpoint of valence one and an unlabelled endpoint of higher valence, then $G/e$ and $G\setminus e$ are homeomorphic, but $G/e$ has one more labelled vertex than $G\setminus e$ (namely the one coming from the higher valence endpoint of $e$ in $G$). See Figure 1.
$G$$e$$G/e$$G\setminus e$

### v2

@@ Abstract.
We show that the complex of partial bases of the free group of rank $n$, where vertices are seen up to conjugation, is Cohen–Macaulay of dimension n1n-1. This confirms a conjecture by Day and Putman. We prove our results in the more general context of freely decomposable groups.
@@ 1. Introduction
1.1. Motivation
Showing that a complex that a group acts on is highly connected (i.e. has vanishing homotopy groups in low degrees) often allows one to deduce algebraic and cohomological properties about the corresponding group. In the present article, the acting group will mostly be the automorphism group of a free group or free product. We begin, however, by discussing the analogous situation for $\operatorname{SL}_{n}(\mathbb{Z})$, which is already better understood. In Section˜1.5 we provide a brief overview of further related settings. For now, we will avoid technical definitions and focus on the bigger picture.
[…]
In the setting of $\operatorname{Out}(F_{n})$, the outer automorphism group of the free group $F_{n}$ of rank $n$, the situation is far less well understood. Nevertheless, there is a commonly used analogue of the symmetric space, namely, the Culler–Vogtmann Outer space [10]. Its simplicial completion, the free splitting complex, can be seen as a bordification of it, which is contractible by [15]. The free splitting complex has several “geometric” descriptions, via marked graphs (of groups) [10], actions on trees [12], and sphere systems in a double handlebody [15]. There is also an “algebraic” description of its boundary: in [6], Brück–Gupta showed that such a boundary has the homotopy type of the complex of free factor systems, defined by Handel–Mosher [13]. They also gave a lower bound on the connectivity of this complex [6, Theorems B and D] and showed that an important subcomplex, the free factor complex, is homotopy equivalent to a wedge of (n2)(n-2)-spheres [^1] An $\operatorname{Aut}(F_{n})$-version of the free factor complex was defined by Hatcher–Vogtmann [16]. It is also known to be (n2)(n-2)-spherical [16], but its homology does not give the dualising module of $\operatorname{Aut}(F_{n})$, as was shown by Himes–Miller–Nariman–Putman [17]. However, this is a proper subcomplex of the complex of free factor systems, and it is still unknown whether the entire boundary is highly connected, i.e. it has the homotopy type of a wedge of spheres of the maximal possible dimension (see [6, Question 1.1] and [25, Question on p. 9]). [^2] Note that the relation between the homology of this boundary complex and the analogue of the Steinberg module is less straightforward in this setting, since Outer space and the free splitting complex are not manifolds. For recent progress in this direction, see the work of Wade–Wasserman [26].
The present article makes a step towards an answer to this question by showing that the complex of free factor systems has many other highly connected subcomplexes (see Theorem˜1.2 and Theorem˜1.3). These results are not sufficient to determine its entire homotopy type, but they let us prove a conjecture by Day–Putman concerning an $\operatorname{Out}(F_{n})$-analogue of the complex of partial bases.
1.2. A conjecture by Day–Putman
Write $[x]$ for the conjugacy class of an element or subgroup $x$ of $F_{n}$. The complex of partial bases of $F_{n}$ is the complex $\mathcal{B}_{n}$ whose maximal simplices are sets $\{C_{1},\ldots,C_{n}\}$ of conjugacy classes such that there exists a basis $\{v_{1},\ldots,v_{n}\}$ of $F_{n}$ with $C_{i}=[v_{i}]$. It is not hard to show that $\mathcal{B}_{n}$ is a pure simplicial complex of dimension n1n-1. This complex was studied by Day and Putman in [11] when looking for generating sets of certain groups of automorphisms of $F_{n}$. Concretely, they provided a topological proof of a theorem by Magnus that describes a generating set for $\mathrm{IA}_{n}$, the Torelli subgroup of $\operatorname{Aut}(F_{n})$. This group is defined as the kernel of the map $\operatorname{Aut}(F_{n})\to\operatorname{GL}_{n}(\mathbb{Z})$. To obtain their result, Day–Putman showed that $\mathcal{B}_{n}$ is connected for $n\geq 2$, simply connected if $n\geq 3$, and that the orbit complex $\mathcal{B}_{n}/\mathrm{IA}_{n}$ is (n2)(n-2)-connected, i.e. spherical of dimension n1n-1. Then they conjecture [11, Conjecture 1.1] that $\mathcal{B}_{n}$ is (n2)(n-2)-connected as well. In this text, we prove their conjecture:
@@ 1.2. A conjecture by Day–Putman
The simplicial complex $\mathcal{B}_{n}$ is Cohen–Macaulay of dimension n1n-1. In particular, $\mathcal{B}_{n}$ is (n2)(n-2)-connected.
1.3. Free factor systems
To prove the above result, we build on and extend work by the first-named author and Gupta [6] and generalisations in [5], which relate $\mathcal{B}_{n}$ to the simplicial boundary of Culler–Vogtmann Outer space [10]. In the present text, we will mostly refer to [4], which contains the material of both [5] and [6]. In fact, we prove a more general version of Theorem˜1.1 that not only applies to free groups but also to free products. To state it, we need to introduce some notation.
Let $A$ be a finitely generated group that is freely decomposable, i.e. is isomorphic to a non-trivial free product. A free factor system of $A$ is a finite set $\mathcal{A}=\left\{[A_{1}],\ldots,[A_{k}]\right\}$ of conjugacy classes of non-trivial proper subgroups $A_{i}\leq A$ such that $A=H\ast A_{1}\ast\cdots\ast A_{k}$, where $H$ is some free subgroup of $A$ and $\ast$ denotes the free product. We allow $\mathcal{A}$ to be empty, but note that the empty set is only a free factor system if $A$ is free itself. The corank of $\mathcal{A}$ is the rank of a free complement $H$, that is, $\operatorname{crk}(\mathcal{A})=\operatorname{rk}(H)$, which is well-defined by the Grushko Theorem.
@@ 1.3. Free factor systems
For $A=F_{n}$ and $\mathcal{A}=\emptyset$, we have c=n2c=n-2 and the result is [6, Theorem B]. In [4, Proposition 4.30], it was shown that $\mathcal{FF}(A,\mathcal{A})$ is homotopy equivalent to the simplicial boundary of relative versions of Culler–Vogtmann Outer space [10] as defined by Guirardel–Levitt [12], so Theorem˜1.2 implies that these boundaries are highly connected.
Note that Theorem˜1.2 does not completely determine the homotopy type of $\mathcal{FF}(A,\mathcal{A})$ as the latter has dimension $2\operatorname{crk}(\mathcal{A})+|\mathcal{A}|-3$, which is greater than $c+1$ in general (see also [4, Question 4.50] and the comments around it). We hope that the strategy presented here and in [6] can be further developed to shed more light on the homotopy type of $\mathcal{FF}(A,\mathcal{A})$.
1.4. Frame complexes
To obtain Theorem˜1.1, we consider a certain subposet of $\mathcal{FF}(A,\mathcal{A})$: The frame complex associated with $(A,\mathcal{A})$, denoted by $\mathcal{PB}(A,\mathcal{A})$, is the subposet of $\mathcal{FF}(A,\mathcal{A})\cup\{[A]\}$ given by free factor systems $\mathcal{A}^{\prime}$ such that for each $A^{\prime}]\in\mathcal{A}^{\prime}$, we either have $A^{\prime}]\in\mathcal{A}$ or $A^{\prime}\cong\mathbb{Z}$. Note that $\mathcal{PB}(A,\mathcal{A})$ is a subposet of $\mathcal{FF}(A,\mathcal{A})$ except if $\mathcal{A}=\emptyset$ and $A\cong\mathbb{Z}$, where $\mathcal{PB}(A,\mathcal{A})=\{[A]\}$. In general, the elements of $\mathcal{PB}(A,\mathcal{A})$ are of the form $\left\{[\langle v_{1}\rangle],\ldots,[\langle v_{i}\rangle]\right\}\sqcup\mathcal{A}$, where $\{v_{1},\ldots,v_{i}\}$ is a partial basis of the free group $H$ in a decomposition $A=H\ast A_{1}\ast\cdots\ast A_{k}$. The poset relation of $\mathcal{FF}(A,\mathcal{A})$ restricts to the subset relation on $\mathcal{PB}(A,\mathcal{A})$, so we can see $\mathcal{PB}(A,\mathcal{A})$ as a simplicial complex with 0-simplices of the form $\left\{[\langle v\rangle]\right\}\sqcup\mathcal{A}$. Since $\mathcal{A}$ is present in any simplex, we could remove it from all of them and get an abstract simplicial complex; however, for consistency, we will keep $\mathcal{A}$ in our simplices and rather think of $\mathcal{PB}(A,\mathcal{A})$ as a poset.
@@ 1.4. Frame complexes
Relation between the theorems
We deduce Theorem˜1.3 from Theorem˜1.2 by showing that the inclusion map $\mathcal{PB}(A,\mathcal{A})\to\mathcal{FF}(A,\mathcal{A})$ is highly connected (see Section 2.1 for definitions). For $A=F_{n}$ and $\mathcal{A}=\emptyset$, we have $\operatorname{crk}(\mathcal{A})=n$ and $\mathcal{PB}(F_{n},\emptyset)$ is an inflation of the complex $\mathcal{B}_{n}$ (in the sense of Definition˜3.5), so Theorem˜1.1 is an immediate consequence of Theorem˜1.3.
1.5. Related work
The complex $\mathcal{B}_{n}$ comes with a natural action of $\operatorname{Out}(F_{n})$, the outer automorphism group of the free group. More generally, relative outer automorphism groups $\operatorname{Out}(A,\mathcal{A}^{t})$ (see [4, Section 4.1.2]) act on $\mathcal{PB}(A,\mathcal{A})$ and $\mathcal{FF}(A,\mathcal{A})$. We would like to mention that variants of the complexes $\mathcal{B}_{n}$ and $\mathcal{PB}(A,\mathcal{A})$ related to several other groups are known to be highly connected as well:
The $\operatorname{Aut}(F_{n})$-version of Theorem˜1.1 was established by Sadofschi Costa [23], who showed that the simplicial complex whose simplices are the partial bases of $F_{n}$ (without modding out by the conjugation action on vertices) is Cohen–Macaulay of dimension n1n-1. In that case, the spanning map yields a poset map from the face poset of the partial basis complex to the free factor poset, which the author leverages via a fibre-type argument relying on the Cohen–Macaulayness of the free factor poset, established by Hatcher–Vogtmann [16]. In our setting, where we work with conjugacy classes, there is no such spanning map. This introduces an additional challenge, which we address by considering product-type configurations and graph posets instead. In particular, our approach differs in technique and is independent of [23].
Furthermore, Maazen [19] showed that a $\operatorname{GL}_{n}$-version of $\mathcal{B}_{n}$ is Cohen–Macaulay; Harer [14] proved that a version related to the mapping class group of a surface is highly connected (cf. the introduction of [11]); Putman studied a version related to the integral symplectic group $\operatorname{Sp}_{2n}(\mathbb{Z})$ [7, 21]. For an overview of how the connectivity of these complexes can be used to compute high-degree cohomology of the corresponding groups, see [8].
@@ Acknowledgements
We thank the anonymous referees for comments that helped to improve the exposition of this article. The first author was supported by the Deutsche Forschungsgemeinschaft (Project 427320536–SFB 1442) and Germany’s Excellence Strategy grant EXC 2044–390685587. The second author was supported by the Alexander von Humboldt Stiftung, and the FWO grant 12K1223N.
@@ 2. Posets of graphs
For the proof of Theorem˜1.2 in Section˜3, we compare $\mathcal{FF}(A,\mathcal{A})$ with a poset of free splittings of $A$, and study connectivity of fibres. These fibres coincide with certain finite posets of graphs, which we show to be highly connected in the present section.
@@ 2.1. Posets and wedges of spheres
We start out by recalling some topological concepts.
@@
Let $m$ be an integer. By convention, every topological space is $m$-connected if $m\leq-2$. A topological space $X$ is (1)(-1)-connected if it is non-empty, and for $m\geq 0$, $X$ is $m$-connected if the homotopy groups of $X$ of degree $\leq m$ vanish regardless of the basepoint. A continuous map $f:X\to Y$ is said to be $m$-connected if it induces an isomorphism between homotopy groups of degree <m<m, and an epimorphism in degree $m$ (again, regardless of the basepoint). If $X,Y$ are finite-dimensional CW-complexes, by a slight abuse of language, we will say that a continuous map $f:X\to Y$ is highly connected if $f$ is an $n$-equivalence, where $n$ is either the dimension of $X$ or of $Y$. When we need to specify the concrete connectivity of the map, we will indicate the grade of the equivalence.
An $n$-dimensional simplicial complex $K$ is spherical if it is (n1)(n-1)-connected. Equivalently, $K$ has the homotopy type of a wedge of $n$-spheres. We say that $K$ is Cohen–Macaulay if for every simplex $\sigma\in K$, its link $\operatorname{lk}_{K}(\sigma)=\{\tau\in K\mathrel{{\>:\>}}\tau\cup\sigma\in K,\tau\cap\sigma=\emptyset\}$ is spherical of dimension $\dim K-\dim\sigma-1$ (and in particular $K$ is spherical by taking $\sigma$ the empty simplex).
If $X$ is a poset, we study its topological properties via its order complex. We call a map $f\colon X\to Y$ between two posets a poset map or an order-preserving map if $x\leq y$ implies $f(x)\leq f(y)$. Any order-preserving map between posets induces a simplicial map, and hence a continuous map, between their order complexes. For two order-preserving maps $f,g:X\to Y$ such that $f\leq g$ (that is, $f(x)\leq g(x)$ for all $x\in X$), the induced maps between the geometric realisations are homotopy equivalent.
If $Y$ is a subposet of $X$, we call $f\colon X\to Y$ monotone if either for all $x\in X$ we have $f(x)\leq x$, or for all $x\in X$ we have $f(x)\geq x$. By [22, 1.3], a monotone poset map $f:X\to X$ defines a homotopy equivalence from $X$ to its image $f(X)$. In particular, a poset with a unique maximal or minimal element is contractible.
We say that $X$ is Cohen–Macaulay if its order complex is. In this case, to check for the Cohen–Macaulay property, one only needs to verify that $X$ is spherical, and that every interval of $X$ is spherical of the right dimension (see, for instance, [22]).
@@ 2.2. Graphs
In what follows, by a graph we mean a $1$-dimensional compact CW-complex with no isolated vertices. For a graph $G$, we denote by $V(G)$ its set of vertices, and by $E(G)$ its set of edges. Note that multiple edges and loops are allowed. By a subgraph we mean an edge-induced subgraph. That is, subgraphs of $G$ are in one-to-one correspondence with subsets of $E(G)$. Recall that the valence of a vertex $v\in V(G)$ is the number of connected components obtained by deleting $v$ from a small enough neighbourhood of $v$, i.e. the number of half-edges adjacent to $v$. An edge that has an endpoint with valence one is called a leaf. An edge is said to be separating if removing any point of its interior disconnects the graph. In particular, we consider leaves to be separating. The rank of a connected graph $G$, denoted by $\operatorname{rk}(G)$, is the rank of its fundamental group.
A labelled graph is a pair $(G,l)$ where $G$ is a graph and $\emptyset\subseteq l\subseteq V(G)$ is a subset of its vertex set. We call $l$ the set of labelled vertices of $G$ and write $k=k(l)=|l|$ for the number of labelled vertices. [^3] We deviate from the notation in [4] here, where the labelling is a map $l:\left\{1,\ldots,k\right\}\to V(G)$. The reason is that we only care about the set of labelled vertices, i.e. the image of such a map $l$ in the present work. A core subgraph of a labelled graph $(G,l)$ is a non-empty subgraph such that every vertex of valence one is labelled. Observe that every labelled graph $(H,l)$ contains a unique maximal core subgraph $\mathring{H}$: this is the subgraph obtained by repeatedly removing leaves with non-labelled endpoints.
Let $e\in E(G)$ be an edge of a labelled graph $(G,l)$. We write $G\setminus e$ for the subgraph of $G$ on $E(G)\setminus\{e\}$ and $G/e$ for the graph obtained from $G$ by collapsing $e$. The graphs $G\setminus e$ and $G/e$ have obvious labellings induced by $l$: the labelling of $G\setminus e$ is given by $l\cap V(G\setminus e)$ and the one of $G/e$ is obtained as the image of $l$ under the quotient map $V(G)\to V(G/e)$. For example, if $e$ is a leaf with a labelled endpoint of valence one and an unlabelled endpoint of higher valence, then $G/e$ and $G\setminus e$ are homeomorphic, but $G/e$ has one more labelled vertex than $G\setminus e$ (namely the one coming from the higher valence endpoint of $e$ in $G$). See Figure˜1.
$G/eGeG\setminus e$